At point P, the angle of elevation of the top of a hill is 37.93°. At point Q on the same horizontal line as P and the foot of the hill and 56.6 meters from P, the angle of elevation is 22.61°. Find the height of the hill.


1
Expert's answer
2021-11-23T08:51:36-0500



Find the length of hypotenuse y

Using sine rule


"\frac{a}{sinA}=\frac{b}{sinB}=\frac{c}{sinC}"


"\frac{y}{sin(22.61\degree)}=\frac{56.6}{sin(15.32\degree)}"


"y=\frac{56.6\times sin(22.61\degree)}{sin(15.32\degree)}"


"y=82.36m"


Now we have the Hypotenuse Y lets find the height h of the hill

"sin\theta=\frac{opposite(h)}{Hypotenuse(y)}"


"h=82.36\times sin(37.93\degree)"


"h=50.63m"


height of the hill is 50.63meters